### v2

# Multivariable Behavioral Change Modeling of Epidemics in the Presence of Undetected Infections

Abstract
Epidemic models are invaluable tools to understand and implement strategies to control the spread of infectious diseases, as well as to inform public health policies and resource allocation. However, current modeling approaches have limitations that reduce their practical utility, such as the exclusion of human behavioral change in response to the epidemic or ignoring the presence of undetected infectious individuals in the population. These limitations became particularly evident during the COVID-19 pandemic, underscoring the need for more accurate and informative models. Motivated by these challenges, we develop a novel Bayesian epidemic modeling framework to better capture the complexities of disease spread by incorporating behavioral responses and undetected infections. In particular, our framework makes three contributions: 1) leveraging additional data on hospitalizations and deaths in modeling the disease dynamics, 2) accounting for data uncertainty arising from the large presence of asymptomatic and undetected infections, and 3) allowing the population behavioral change to be dynamically influenced by multiple data sources (cases and deaths). We thoroughly investigate the properties of the proposed model via simulation, and illustrate its utility on COVID-19 data from Montréal and Miami.
1 Division of Biostatistics and Health Data Science, University of Minnesota
2 Faculty of Veterinary Medicine, University of Calgary
3 Department of Mathematics and Statistics, University of Calgary
4 Department of Epidemiology, Biostatistics, and Occupational Health, McGill University
Keywords Bayesian inference $\cdot$ Compartmental model $\cdot$ Transmission model $\cdot$ SIHRD $\cdot$ SIR
1 Introduction
Epidemic models are essential tools for understanding transmission dynamics and informing public health responses. The foundational SIR compartmental model (Kermack and McKendrick,, 1927), describes disease dynamics by classifying individuals as Susceptible, Infectious, or Removed and using parameters to describe the rates of flow between compartments. These models can be solved deterministically with ordinary differential equations, but stochastic Bayesian implementations are often preferred due to their ability to incorporate prior knowledge and impute missing data through data augmentation (O’Neill and Roberts,, 1999; Lekone and Finkenstädt,, 2006). Most SIR models neglect the role of human behavior changes in shaping transmission dynamics, however, this is a critical limitation that reduces their practical utility and accuracy, a fact that was exposed by the COVID-19 pandemic.
Several behavioral change SIR models have been proposed in the deterministic literature (see Funk et al., (2010) for a comprehensive review), however, these deterministic approaches do not perform parameter estimation using observed data. Recently a stochastic Bayesian behavioral change modeling framework was developed where transmission is dynamically modified by a so-called ‘alarm’ function, which captures the time-varying changes in the population engaging in protective behaviors Ward et al., (2023, 2025). This framework allows for inference on transmission and behavioral parameters and was shown to provide superior model fit to outbreaks of COVID-19, Ebola, and Foot and Mouth Disease, compared to models ignoring behavioral change. Despite this success, these models rely on simplifying assumptions: the use of a standard SIR structure and a behavioral alarm that depends on only one measure of disease severity (incidence or prevalence).
Given the increased availability of hospitalization and death data during an outbreak, there is strong motivation to develop data-driven models that account for these disease states. However, most fully stochastic Bayesian epidemic models are limited to simple SIR or SEIR compartmental structures to reduce computational complexity (Andersson and Britton,, 2012). While expanded compartmental structures (e.g., SIHVR to include hospitalizations and vaccinations (Gibson et al.,, 2023)) have been explored in computationally efficient, partially stochastic models, these approaches neglect the important uncertainty in unobserved compartments by modeling transitions deterministically.
Throughout the COVID-19 pandemic, modeling has been complicated by the large presence of undetected infections (Menachemi,, 2020; Irons and Raftery,, 2021). Existing deterministic or partially stochastic models have addressed this by adding an undetected compartment (Melis and Littera,, 2021; Bhaduri et al.,, 2022) or using a multiplier on the observed case counts (Hao et al.,, 2020; Ma and Rennert,, 2024). In fully stochastic individual-level models, undetected infections can be estimated using reversible jump Markov Chain Monte Carlo (MCMC) techniques (Forrester et al.,, 2007; Jewell and Roberts,, 2012; Ward et al.,, 2021). However, these approaches are computationally infeasible for population-averaged models in large populations as they require imputing disease status for each individual. Thus, there is a need to develop efficient methodology for incorporating undetected infections in fully stochastic population-averaged models.
Motivated by these challenges, we extend the fully stochastic population-averaged behavioral change model to 1) include hospitalizations and deaths in an SIHRD model, 2) incorporate undetected infections, and 3) use a multivariable alarm to quantify the relative importance of two data sources on behavior. In Section 2, we review the univariable alarm SIR model and present our extended SIHRD model with a multivariable alarm and incorporating undetected infections. In Section 3, the proposed methodology is rigorously investigated through simulations. An analysis of two waves of COVID-19 in Miami, Florida and Montréal, Québec is performed in Section 4 to illustrate the utility and limitations of the proposed modeling approach on real data. Concluding remarks are provided in Section 5.
2 Methods
2.1 Behavioral Change SIR Model
We start by briefly introducing the discrete time behavioral change SIR model of Ward et al., (2023) which serves as the foundation for our work. Assume a closed population of size $N$ and let $S_{t}$, $I_{t}$, and $R_{t}$ denote the number of individuals in the susceptible, infectious, and removed compartments in the continuous time interval $[t,t+1)$, respectively. Transition vectors $I^{*}_{t}$ and $R^{*}_{t}$ represent the number of individuals entering the indicated compartment in this interval. Transitions are temporally described by
$S_{t+1}$ $=S_{t}-I^{*}_{t}$
$I_{t+1}$ $=I_{t}+I^{*}_{t}-{R_{t}}^{*}$
$R_{t+1}$ $=R_{t}+{R_{t}}^{*}.$
Given counts in each compartment at time 0 and the transition vectors, $\boldsymbol{S}$, $\boldsymbol{I}$, and $\boldsymbol{R}$ are fully determined by these difference equations. The initial states can be fixed or estimated using a multinomial prior. Transitions between compartments are binomially distributed as $I^{*}_{t}\sim Bin\left(S_{t},\pi_{t}^{(SI)}\right)$ and $R^{*}_{t}\sim Bin\left(I_{t},\pi^{(IR)}\right)$.
Of primary interest is the transmission probability $\pi_{t}^{(SI)}$, traditionally specified as $\pi_{t}^{(SI)}=1-\exp\left(-\beta\frac{I_{t}}{N}\right)$. This probability is derived from assumptions of an independent Poisson contact process and constant infection probability given a contact, with $\beta$ capturing both the contact rate and infection probability as these are not separately identifiable (Brown et al.,, 2016). Behavioral change is incorporated by allowing the transmission rate, $\beta$, to be modified by a time-varying level of alarm, $a_{t}$, in the population as
$\pi_{t}^{(SI)}=1-\exp\left\{-\beta(1-a_{t})\frac{I_{t}}{N}\right\}.$ (1)
The alarm is constrained to be in the interval $[0,1]$, and is interpreted as the proportional reduction in transmission due to the alarm in the population. When $a_{t}=0$, the population is in its natural ‘unalarmed’ state, and transmission is described only by $\beta$. When $a_{t}=1$, the population is in its maximal alarmed state and transmission is reduced to zero. In Ward et al., (2023), the alarm was specified as a function of previously observed incidence smoothed over the past $m$ days, i.e., $a_{t}=f\left(I^{*}_{t-1,m}\right)$, where $I^{*}_{t-1,m}=\frac{1}{m}\sum_{i=t-m-1}^{t-1}I^{*}_{i}$ with smoothing parameter $m\in\{1,2,...,t-1\}$. Several functions have been considered to describe the alarm, and we will use the one parameter ‘power’ alarm with growth rate $k>0$
$a_{t}=1-(1-I^{*}_{t-1,m}/N)^{1/k}.$ (2)
The removal probability is typically specified as $\pi^{(IR)}=1-\exp\left(-\gamma\right)$, which derives from an exponentially distributed length of the infectious period. However, this may be a restrictive assumption and without complete data on both $\boldsymbol{I^{*}}$ and $\boldsymbol{R^{*}}$ would require data-augmented MCMC approaches to implement (Lekone and Finkenstädt,, 2006). To incorporate more flexible modeling of the infectious period and avoid data augmentation, we use the infectious duration-dependent (IDD) specification of Ward et al., (2022), which was shown to be computationally efficient and to improve estimation relative to traditional data augmentation. Let $I_{wt}$ be the number of individuals on day $w$ of the infectious period at time $t$ and $T_{I}$ denote the fixed length of the infectious period. The IDD transmission probability is
$\pi_{t}^{(SI)}=1-\exp\left\{-\beta(1-a_{t})\frac{\sum_{w=1}^{T_{I}}f(w)I_{wt}}{N}\right\},$ (3)
where $f(w)$ is a function defining the weighted contribution of infectious individuals depending on how long they’ve been infectious. In our implementation, we use a logistic decay function with a maximum curve value of 1, $f(w)=1/[1+\exp\{\nu(w-w_{0})\}]$, where $w_{0}$ provides the inflection point of the curve, and $\nu>0$ provides the decay rate. As the length of the infectious period is fixed, removal times are fully specified as $R^{*}_{t+T_{I}}=I^{*}_{t}$, for all $t$ and we no longer need to estimate the removal probability $\pi^{(IR)}$. Instead, the parameters of the logistic decay curve are used to quantify the length of time infectious individuals are transmitting the disease. The logistic decay curve assumes peak transmissibility at the start of the infectious period, which may not always be realistic. Alternative IDD curves can allow later peaks, but Ward et al., (2022) found minimal changes in overall transmission estimation between various curves and the logistic decay curve was found to have the highest MCMC efficiency.
2.2 Multivariable Alarm in an SIHRD model with Undetected Infections
We extend the behavioral change SIR model in three important ways. First, we consider an expanded SIHRD compartmental structure to include hospitalizations (H) and deaths (D). The SIHRD model is temporally described by the following set of difference equations:
$S_{t+1}$ $=$ $S_{t}-I^{*}_{t}$
$I_{t+1}$ $=$ $I_{t}+I^{*}_{t}-H^{*}_{t}-{R_{t}^{I}}^{*}$
$H_{t+1}$ $=$ $H_{t}+H^{*}_{t}-{R_{t}^{H}}^{*}-D^{*}_{t}$
$R_{t+1}$ $=$ $R_{t}+{R_{t}^{I}}^{*}+{R_{t}^{H}}^{*}$
$D_{t+1}$ $=$ $D_{t}+D^{*}_{t},$
where $I_{t}^{*}$, $H_{t}^{*}$, and $D_{t}^{*}$ correspond with the observed incidence, new hospitalizations, and new deaths over time. $R_{t}^{I}$$^{*}$ and $R_{t}^{H}$$^{*}$ denote the number of new recoveries from the I and H compartments, respectively, which are unobserved. As written, these equations imply that infectious individuals may become hospitalized or may recover without hospitalization, and that any deaths due to the disease must first be hospitalized. While it would be possible to rewrite the equations such that infectious individuals are allowed to die without hospitalization, this introduces additional complexity to the model which is not likely to be supported by publicly available data (i.e., $D_{t}^{*}$ is observed, but $D_{t}^{I}$$^{*}$ and $D_{t}^{H}$$^{*}$ are not). Transitions between compartments are probabilistically described as
$I_{t}^{*}$ $\sim Binom\left(S_{t},\pi_{t}^{(SI)}\right)$
$\{H_{t}^{*},{R_{t}^{I}}^{*},I_{t+1}\}$ $\sim Multinom\left(I_{t},\left\{\pi^{(IH)},\pi^{(IR)},1-\pi^{(IH)}-\pi^{(IR)}\right\}\right)$
$\{D_{t}^{*},{R_{t}^{H}}^{*},H_{t+1}\}$ $\sim Multinom\left(H_{t},\left\{\pi^{(HD)},\pi^{(HR)},1-\pi^{(HD)}-\pi^{(HR)}\right\}\right).$
The multinomial distribution is used to describe transitions from I and H as a straightforward extension of the original chain binomial SIR model. Infectious individuals may become hospitalized, recover, or stay infectious according to the probabilities $\pi^{(IH)}$, $\pi^{(IR)}$, or $1-\pi^{(IH)}-\pi^{(IR)}$, respectively. Similarly, hospitalized individuals may die, recover, or remain hospitalized. We assume exponentially distributed lengths of time in non-susceptible compartments, leading to the following transition probabilities: $\pi^{(IH)}=1-\exp\left(-\lambda\right)$, $\pi^{(IR)}=1-\exp\left(-\gamma_{1}\right)$, $\pi^{(HR)}=1-\exp\left(-\gamma_{2}\right)$, $\pi^{(HD)}=1-\exp\left(-\phi\right)$.
The second extension to the model is to allow for the presence of undetected infections. This captures any individual who did not have a confirmed positive test reported to health officials. Undetected infections are incorporated by splitting the infectious compartment into two parts: observed (detected) cases, $C_{t}$, and undetected infectious individuals, $U_{t}$, such that $I_{t}=C_{t}+U_{t}$. Correspondingly, $I_{t}^{*}=C_{t}^{*}+U_{t}^{*}$, where $C_{t}^{*}$ are the observed cases counts and $U_{t}^{*}$ is unobserved. We assume $C_{t}^{*}\sim Binom\left(I_{t}^{*},\pi^{detect}\right)$, where $\pi^{detect}$ is the probability of an infection being detected. In many cases, there may be data from seroprevalence studies available to inform a strong prior on this detection probability. The proposed SIHRD model with undetected infections is visualized in Web Appendix A.
Lastly, we consider extending the alarm function of Ward et al., (2023) to allow multiple data sources to inform population behavior. The transmission probability $\pi_{t}^{(SI)}$ remains unchanged from Equation 1, but we now consider the alarm to be a function of observed cases and deaths, i.e., $a_{t}=f(C_{t-1,m}^{*},D^{*}_{t-1,m})$. While theoretically the alarm could also depend on other factors, we found that when several highly correlated data sources were used (e.g., cases, hospitalizations, and deaths), a lack of identifiability inhibited estimation. Cases and deaths were chosen as those are most commonly available and because they are furthest apart in the disease process which reduces correlation. In a model with undetected infections the alarm could be specified as a function of the true incidence, $I_{t-1,m}^{*}$, however, as this is generally unknown to the population during an outbreak we find it more realistic to assume that behavior is only informed by observed data. We define the multivariable alarm as
$a_{t}=1-\left[1-\left\{\alpha C_{t-1,m}^{*}+(1-\alpha)D^{*}_{t-1,m}\right\}/N\right]^{1/k}.$ (4)
This parameterization was chosen for interpretability of the $\alpha$ parameter as the relative contribution of observed cases in informing the population alarm compared to deaths. Thus, the multivariable alarm has advantages of being more flexible than a single metric alarm while also providing inference to help understand what influences population behavior.
2.3 Reproductive Number Calculation
Often, the primary inferential interest in compartmental epidemic models is the reproductive number as it quantifies the magnitude of disease spread in a population. We consider the time-varying “effective” reproductive number, $\mathcal{R}_{0}(t)$, defined as the expected number of secondary infections caused by a single infected individual in the current population at time $t$, as our interest is in incorporating behavioral change leading to changing transmission over time. As shown in Ward et al., (2022), this can be generally derived as an expectation from the chain binomial SIR model presented in Section 2.1 as $\mathcal{R}_{0}(t)=\sum_{j=t}^{\infty}S_{j}\pi_{0j}^{(SI)}\left(1-\pi^{(IR)}\right)^{j-t}$, where for a single individual infected at time $t$ $\pi_{0j}^{(SI)}$ denotes the probability of transmission from that individual at time $j$. Intuitively, the formulation is derived from assuming that on any given day, the expected number of secondary infections is $S_{j}\times\pi_{0j}^{(SI)}$, and then over the course of the infection this quantity is weighted by the probability that the individual is still infectious (i.e., has not become removed). For the IDD transmissibility specification we are using to implement the SIR model, the duration of the infection is fixed so this simplifies to $\mathcal{R}_{0}(t)=\sum_{j=t}^{t+T_{I}}S_{j}\Big[1-\exp\Big\{-\beta(1-a_{j})\frac{f(j-t)}{N}\Big\}\Big]$. This formulation can be directly extended to the SIHRD compartmental structure, however, we must be sure to account for the multinomial transition out of the Infectious compartment as according to the model individuals can stop transmitting the disease when they recover or when they become hospitalized. Thus, the effective reproductive number is calculated as $\mathcal{R}_{0}(t)=\sum_{j=t}^{\infty}S_{j}\Big[1-\exp\Big\{-\beta(1-a_{j})\frac{1}{N}\Big\}\Big]\left(1-\pi^{(IH)}-\pi^{(IR)}\right)^{j-t}$.
2.4 Estimation
Inference is performed in the Bayesian framework. The log-likelihood for the SIR model with IDD transmission is $\ell(\mathbf{I}^{*},\mathbf{R}^{*}|\boldsymbol{\Theta})=\sum_{t=0}^{\tau}\left\{\log\binom{S_{t}}{I^{*}_{t}}+I^{*}_{t}\log\pi_{t}^{(SI)}+\left(S_{t}-I^{*}_{t}\right)\log\left(1-\pi_{t}^{(SI)}\right)\right\}$, with parameter vector $\Theta$ containing $\beta$, $w_{0}$, $\nu$ and any parameters associated with an alarm function (e.g., $k$ and $\alpha$ for the multivariable alarm). Commonly, $\beta$ and $k$ are given vague gamma priors (e.g., Gamma(0.1, 0.1)), and informative priors are used for the parameters $w_{0}$ and $\nu$ which describe the logistic decay curve of transmissibility over an individual’s infection. As $w_{0}$ is the inflection point it can be thought of as the average length of time individuals are infectious using a normal prior with small variance. For the decay rate, $\nu$, we typically use Gamma priors centered on a value of 1 (e.g., Gamma(100, 100)) to enforce a sharp decrease in transmission around the inflection point for interpretability of $w_{0}$.
For the proposed SIHRD model with undetected infections the complete log-likelihood is
$\ell(\mathbf{I}^{*},\mathbf{C}^{*},\mathbf{H}^{*},\mathbf{R^{I}}^{*},\mathbf{R^{H}}^{*},\mathbf{D}^{*}|\boldsymbol{\Theta})=$
$\hskip 25.60747pt\sum_{t=0}^{\tau}\Big\{\log\binom{S_{t}}{I^{*}_{t}}+I^{*}_{t}\log\pi_{t}^{(SI)}+\left(S_{t}-I^{*}_{t}\right)\log\left(1-\pi_{t}^{(SI)}\right)$
$\hskip 25.60747pt+\log\frac{I_{t}!}{H^{*}_{t}!{R^{I}_{t}}^{*}!I_{t+1}!}+H^{*}_{t}\log\pi^{(IH)}+{R^{I}_{t}}^{*}\log\pi^{(IR)}+I_{t+1}\log\left(1-\pi^{(IH)}-\pi^{(IR)}\right)$
$\hskip 25.60747pt+\log\frac{H_{t}!}{D^{*}_{t}!{R^{H}_{t}}^{*}!H_{t+1}!}+D^{*}_{t}\log\pi^{(HD)}+{R^{H}_{t}}^{*}\log\pi^{(HR)}+H_{t+1}\log\left(1-\pi^{(HD)}-\pi^{(HR)}\right)$
$\hskip 25.60747pt+\log\binom{I^{*}_{t}}{C^{*}_{t}}+C^{*}_{t}\log\pi^{detect}+\left(I^{*}_{t}-C^{*}_{t}\right)\log\left(1-\pi^{detect}\right)\Big\},$
with parameter vector $\Theta$ containing $\beta$, $\gamma_{1}$, $\gamma_{2}$, $\lambda$, $\phi$, $\pi^{detect}$, $k$ and $\alpha$ for the model with the multivariable alarm function. Complete data would provide $\mathbf{I}^{*}$, $\mathbf{C}^{*}$, $\mathbf{H}^{*}$, $\mathbf{R^{I}}^{*}$, $\mathbf{R^{H}}^{*}$, and $\mathbf{D}^{*}$ and initial conditions ($S_{0}$, $I_{0}$, $H_{0}$, $R_{0}$, $D_{0}$), however, from publicly available data we typically only observe $\mathbf{C}^{*}$, $\mathbf{D}^{*}$, and sometimes $\mathbf{H}^{*}$. We assume these three data sources are available and use the data-augmented MCMC approach of Lekone and Finkenstädt, (2006) to impute $\mathbf{R^{I}}^{*}$ and $\mathbf{R^{H}}^{*}$. We are not able to avoid their imputation through the IDD approach taken in the SIR model due to the multinomial transitions from the I and H compartments as the IDD specification has not yet been extended for this purpose. Data augmentation is also used to estimate the unobserved true incidence $\mathbf{I}^{*}$ in models that incorporate undetected infections. When undetected infections are ignored it is assumed that $\mathbf{I}^{*}=\mathbf{C}^{*}$ and the last component of the likelihood above is not included.
As in the SIR model, vague Gamma priors are used for $\beta$ and $k$ in the SIHRD model. Informative Gamma priors may be used for the parameters $\gamma_{1}$ (I to R), $\gamma_{2}$ (H to R), $\lambda$ (I to H), and $\phi$ (H to D) in the presence of prior knowledge about the disease state. For imputing $\mathbf{I}^{*}$, strong priors are required for $\gamma_{1}$ and $\lambda$ to induce identifiability. Otherwise, the same overall reproductive number could be produced from a high transmission rate and short infectious periods or a relatively lower transmission rate with long infectious periods. When hospitalizations and deaths are observed, less informative priors can be used for $\lambda$ and $\phi$ as data is available to estimate these parameters. However, more informative priors on $\gamma_{1}$ and $\gamma_{2}$ can greatly aid in estimation and are often available based on the interpretation of $1/\gamma_{1}$ ($1/\gamma_{2}$) as the average length of time spent infectious (hospitalized) before recovery. We assign $\pi^{detect}$ a strong Beta prior centered on the mean used to define $\mathbf{I}^{*}$. The relative importance parameter $\alpha$ is given a Uniform(0, 1) prior as our interest lies in estimating this from observed data without any strong prior assumptions.
All models were implemented using the R package nimble (de Valpine et al.,, 2017, 2021) with user-defined samplers for the data-augmented MCMC algorithms. For all analyses, three MCMC chains were run using various starting values of the parameters and convergence was assessed by ensuring the Gelman and Rubin diagnostic value for all parameters fell below 1.1 (Gelman et al.,, 1992). All code to reproduce the simulations and data analyses presented is available at https://github.com/ceward18/multipleDataBCM.
3 Simulation Study
3.1 Simulation Set-up
A simulation study was designed to achieve the following objectives: 1) how estimation is impacted if the presence of undetected cases is ignored, 2) whether the full SIHRD model is necessary or if a simpler SIR model can still provide the same inferential results, and 3) if the multivariable alarm is estimable and how it affects model fit compared to a univariate alarm or ignoring behavioral change completely. To achieve these goals, epidemics were simulated under the full SIHRD model with undetected infections and alarm based on smoothed cases and deaths as specified in Equation 4 with smoothing over the past $m=30$ days. All epidemics were simulated in a population of one million people with five initially infected individuals and were recorded for 40 days. The detection probability was set to $\pi^{detect}=$ 25% for all simulations. Three values of the weight parameter $\alpha$ in the alarm function were used for data generation: $\alpha=0.85$ for high case importance, $\alpha=0.5$ for equal importance of cases and deaths, and $\alpha=0.15$ for high death importance. 50 epidemics were simulated in each of the three data generating scenario. All other model parameters were given the same value across the three scenarios (full specifications are provided in Web Appendix A). This led to epidemics as illustrated in Figure 1.
Figure 1: Simulated epidemics from three scenarios capturing various importance of cases and deaths informing behavioral change. Plotted over time are the total (unobserved) infections (dashed gray), observed cases (solid black), hospitalizations (solid gold), and deaths (solid blue). In the high deaths importance scenario, infections peak at counts around 3,000 but this was excluded from the figure to highlight more subtle differences between high case importance and equal importance generating scenarios.
Six models were fitted to each simulated epidemic to enumerate all possible combinations of the two compartmental structures (SIHRD and SIR) and three alarm specifications: multivariable alarm based on cases and deaths, alarm based on cases only as specified in Equation 2, and no alarm ($a_{t}=0$ for all $t$). Each of these models was fit under two assumptions: observed cases is the true incidence (ignoring undetected infections) and allowing for undetected infections to be incorporated in the modeling process as specified in Section 2.4. When undetected infections were modeled, we assume a correctly specified prior, such that $E(\pi^{detect})=0.25$ with small variance. Other priors were specified as in Section 2.4, and a full description is provided in the Web Appendix A. Three MCMC chains with different initial values were run for each model with a burn-in of 100,000 iterations and 300,000 post burn-in samples drawn every 10th iteration. Convergence was established by a Gelman and Rubin diagnostic value below 1.1 (Gelman et al.,, 1992). A small number of the SIHRD models (<< 1%) did not converge and have been excluded from the results.
3.2 Simulation Results
We use estimation of the reproductive number to assess the impact of ignoring undetected cases and to compare the SIHRD and SIR compartmental structures due to its epidemiological importance. Across simulations k=1,,Kk=1,...,K, each model’s estimated $\hat{\mathcal{R}}_{0}(t)$ was computed and compared to the true $\mathcal{R}_{0}(t)$ using the root mean squared error (RMSE) computed as $\sqrt{\frac{1}{K}\sum_{k=1}^{K}\left\{\hat{\mathcal{R}}_{0}(t)-\mathcal{R}_{0}(t)\right\}^{2}}$ at each time point $t$. In Figure 2, we present RMSE at the first and last time point to illustrate estimation prior to behavioral change occurring (start) and after behavioral change has reduced transmission (end). RMSE and the posterior distribution of $\mathcal{R}_{0}(t)$ from one simulation across all time points are provided in Web Appendix A.
Figure 2: Root mean squared error (RMSE) of the effective reproductive number, $\mathcal{R}_{0}(t)$, at the start and end of the simulated epidemics for the three data generating scenarios, the SIHRD and SIR models, and modeling or ignoring undetected infections.
First, we find that models that do not incorporate any behavioral change have very poor estimation of transmission at both the start and end of the epidemics, as indicated by the large RMSE values for the models with no alarm. In assessing estimation of the reproductive number when the presence of undetected cases is ignored, we observe only minimal differences in RMSE of $\mathcal{R}_{0}(t)$, with the most improvement in RMSE seen for the SIHRD models. This result is somewhat unexpected as the true number of infections is four times higher than the observed case counts and ignoring these seems like it should negatively bias the reproductive number estimates. Inspecting the posterior distributions of all parameters in the SIHRD models (Web Appendix A Figures S4-S6), we find that while the transmission parameters ($\beta$, $k$, $\alpha$) were generally similar in distribution whether or not undetected infections were incorporated into the modeling, the parameters associated with the length of the infectious period were biased such that $\pi^{(IH)}$ was larger and $\pi^{(IR)}$ was smaller than their true values. Since hospitalizations over time are observed, ignoring undetected infections increases $\pi^{(IH)}$ as there are less infected people available to be hospitalized. In combination, the tradeoff between these two probabilities ends up having minimal impact on the estimation of $\mathcal{R}_{0}(t)$. Evaluating the posterior distributions of SIR model parameters (Web Appendix A Figures S7-S9), we see minimal changes in the posterior distributions when undetected infections were incorporated or ignored. Thus any differences in the reproductive number would be a result of differences in the number of susceptibles over time. However, as the simulated epidemics are infecting a relatively low proportion of individuals in a large population, incorporating undetected infections has only a minor impact on the total number of susceptible individuals and therefore on the estimated $\hat{\mathcal{R}}_{0}(t)$. We anticipate the incorporation of undetected infections would offer more improvement for epidemics occurring in small populations.
Next we evaluate Figure 2 for differences in the RMSE of $\mathcal{R}_{0}(t)$ between the two compartmental structures and the full multivariable alarm compared to the reduced alarm based on cases only. For both modeling choices, we find that there is little penalty in using the simpler model or alarm structure in terms of estimation of $\mathcal{R}_{0}(t)$ in the high case importance and equal importance data generating scenarios. Where it becomes most important to have the correct model specification is in the scenario where deaths are highly important in driving behavior and one is interested in estimating $\mathcal{R}_{0}(t)$ towards the end of the epidemic after behavioral changes have occurred. It’s intuitive that a model ignoring deaths (i.e., the cases only model) would do poorly in this scenario where deaths are highly important. It is perhaps more interesting that even when the alarm function was correctly specified to include both cases and deaths, if the SIR compartmental structure was used $\mathcal{R}_{0}$ estimation suffers. This indicates that the more complex compartmental structure provided by the SIHRD model is important to ensure accurate estimation of the reproductive number, even when the transmission rate is correctly modeled.
Finally, to evaluate the estimation of the multivariable alarm we focus on estimation of $\alpha$, the relative importance of cases in the alarm function across both the SIHRD and SIR models for all simulations (Figure 3). These results illustrate that the SIR model is biased towards cases being more important, whereas the SIHRD model does a much better job of capturing this parameter, particularly in the high deaths importance and equal importance scenarios. This is also likely the source of the increase in RMSE of $\mathcal{R}_{0}$ for the SIR model in the high deaths importance scenario observed in Figure 2. We hypothesize that this may be due to the fact that deaths are directly included into the compartmental structure in the SIHRD model. The SIHRD model does still show a small amount of bias towards higher case importance as reflected by point estimates being consistently above the true value. The estimation of $\alpha$ is worst in the equal importance scenario, but this is not necessarily surprising as we see that epidemics generated under high case importance and equal importance scenarios are more similar than epidemics generated under the high deaths importance scenario (Figure 1).
Figure 3: Posterior means and 95% credible intervals for $\alpha$ across all simulations for the three data generating scenarios (columns) and the SIHRD and SIR models with multivariable behavioral change as specified in Equation 4 (rows). The true parameters used in simulation are shown by the dashed red lines.
4 Data Analysis
4.1 COVID-19 in Miami and Montréal
To illustrate the model utility in practice, we analyzed data from the first two waves of COVID-19 in Miami-Dade County (Miami), Florida in the United States and Montréal, Québec in Canada. These locations were chosen as both were severely affected by COVID-19, with Montréal having the highest death rate in Canada and Miami-Dade County having the third highest case rate in the United States (NYT COVID tracker). These two locations also provide an interesting comparative case study as they have very different trends in case and death rates over time. As illustrated by Figure 4, while Miami saw a significant increase in cases and a small increase in deaths in the second wave, Montréal had similar case rates during both waves, but a large decrease in the death rate during the second wave.
Case and death data for Miami-Dade county was obtained from the New York Times, based on reports from state and local health agencies (The New York Times,, 2024). Miami hospitalization data was obtained from the Miami-Dade COVID Project (Williams et al.,, 2021, 2024) and provides the number of COVID-19 positive patients admitted on each day. The case and death counts start on March 11, 2020 when the first case occurred in the county, however hospitalization data is not available until April 2, 2020. The missing hospitalization data was imputed using data augmentation. The population size was set at $N=$ 2,701,767, the value reported by the 2020 U.S. census. Daily case, hospitalization, and death counts for Montréal were obtained from theNational Institute of Public Health of Quebecs, (2024), with the first positive case reported on March 5, 2020. The population size for Montréal was set to $N=$ 1,762,949 as this was the value reported in the 2021 Canadian Census of Population.
Figure 4: COVID-19 case and death rates per 100,000 during 2020 in Miami-Dade County, Florida and Montréal, Québec. Gray shaded regions are those used in model fitting. In Miami, Wave 1 started on March 11, 2020 and Wave 2 started on June 1. In Montréal, Wave 1 started on March 5, 2020 and Wave 2 started on August 23, 2020.
To illustrate the use of these models during an active outbreak, models were fit to the first six to ten weeks of each wave. The amount of data included varied by city and wave, but was chosen to include only a few weeks of data after the peak. Sensitivity analyses using different numbers of weeks of data in model fitting are presented in the Web Appendix A. As in the simulation study, six models were fitted to encompass both compartmental structures and three alarm specifications. Despite the simulation results indicating minimal benefit of incorporating undetected infections, all models incorporated this uncertainty. Prior distributions were specified as in Section 2.4 and are detailed in the Web Appendix A. Priors for the probability of detection in Miami were specified based on the Institute for Health Metrics and Evaluation (IHME) estimates of total infections and reported cases in Florida during the corresponding waves, leading to prior means of 12% in wave 1 and 34% in wave 2 (for Health Metrics and, IHME). For Montréal, these priors were based on IHME COVID-19 policy briefing on the percent of COVID-19 infections detected across Canada, with prior means of 25% detection in wave 1 and 40% detection in wave 2 (Institute for Health Metrics and Evaluation , 2022, IHME). Priors for the removal rate for hospitalized patients were chosen to reflect an average length of stay of 15 days (Alimohamadi et al.,, 2022) and priors for the removal rate and IDD curve for infected but not hospitalized individuals were chosen to reflect an average seven day infectious period before recovery. As in the simulation study, three chains were run and convergence established by a Gelman Rubin diagnostic value below 1.1.
4.2 Results
First, we evaluate model fit and compare the estimates of the relative importance of cases ($\alpha$) in informing the alarm for both cities using the SIHRD and SIR models (Table 1). We use the Widely Applicable Information Criteria (WAIC) to evaluate model fit (Watanabe and Opper,, 2010). WAIC is a likelihood-based estimate of prediction error which includes a penalty for the effective number of model parameters (Gelman et al.,, 2013). Since the SIR and SIHRD models have different likelihoods, the WAIC values are not comparable across the two compartmental structures, but it can still be used to determine the best specification of transmission rate within each structure as shown by (Ward et al.,, 2023).
During the first wave of COVID-19 cases in both cities, the estimated relative importance parameter $\alpha$ in the multivariable alarm model indicates that the majority ($\sim$88-94%) of the population alarm is due to cases during the first wave. This concurs with the WAIC values, which are essentially equivalent between the multivariable alarm and the cases only alarm in both cities. However, in Wave 2 in Miami we found that deaths were highly influential as the estimated $\hat{\alpha}$ values are close to zero and WAIC shows that model fit is significantly improved through the addition of deaths to the alarm function for both the SIHRD and SIR models. In Wave 2 in Montréal, the models estimate relatively equal contribution of case and deaths inform alarm, however the credible intervals are extremely wide. These wide intervals are likely due to the low number of deaths during the time period of model fitting (Figure 4). Again the WAIC values differ by <1<1 point between the cases and deaths alarm and the cases only alarm model. Overall the WAIC values indicate that there is little downside in terms of model fit for incorporating deaths into the alarm function even when they are not very important, and that model fit is much worse when behavioral change is not incorporated at all.
SIHRD SIR
City Wave Model fitted WAIC $\hat{\alpha}$ WAIC $\hat{\alpha}$
Cases + deaths 627.0 0.875 (0.619, 0.996) 516.2 0.881 (0.636, 0.996)
Cases only 625.2 - 515.1 -
Wave 1 No alarm 1723.0 - 1141.6 -
Cases + deaths 1649.1 0.000 (0.000, 0.000) 1132.8 0.002 (0.001, 0.003)
Cases only 2156.0 - 1391.0 -
Miami Wave 2 No alarm 3596.5 - 1420.8 -
Cases + deaths 2454.3 0.941 (0.800, 0.998) 739.3 0.942 (0.805, 0.998)
Cases only 2454.5 - 738.6 -
Wave 1 No alarm 3006.7 - 875.2 -
Cases + deaths 1349.8 0.456 (0.023, 0.972) 576.1 0.616 (0.149, 0.983)
Cases only 1349.3 - 575.6 -
Montréal Wave 2 No alarm 1520.8 - 647.8 -
Table 1: WAIC values for all models and posterior means (95% credible intervals) of $\alpha$ for multivariable alarm models. SIHRD and SIR models have different likelihoods so the WAIC values are not comparable across compartmental structures. Lowest WAIC within each wave/compartment grouping is shaded. The parameter $\alpha$ provides the relative importance of cases in informing the population alarm, with values closer to one indicating high case importance and values closer to zero indicating high death importance.
Evaluating the posterior distributions of the alarm functions and reproductive numbers over time illustrates several other important findings (Figure 5). In both cities, Wave 1 corresponds with a quickly increasing alarm function, reaching an almost 70% reduction in transmission in six weeks in Miami and around 50% reduction in transmission in Montréal after eight weeks. Correspondingly, the reproductive number in both cities, which is estimated to be between 2 and 3 for both cities at the start of the pandemic decreases sharply, although it does not quite cross the threshold of one during the time period of modeling. Models which do not incorporate behavioral change through the alarm function estimate a constant reproductive number which is much lower (around 1.1 - 1.2), as these models do not contain the flexibility to allow transmission to change over time. During the second wave, the alarm functions no longer start at zero as it is not the start of the epidemic and cases and deaths are still occurring at low levels prior to the wave starting. In Miami, the alarm estimation during the second wave is heavily dependent on the inclusion of deaths in the alarm function, as our estimation of $\alpha$ showed deaths were the primary influence of behavioral change during this wave. Deaths were already elevated at the start of the wave, and alarm increases after around 40-50 days once the death rate started rapidly increasing in mid-July. The estimated reproductive number is about 1.25 in Miami and about 1.5 in Montréal at the start of Wave 2. In both cities the behavioral change models do find that $\hat{\mathcal{R}_{0}}(t)$ becomes below 1 by the end of the modeling period. As in Wave 1, the models with no alarm results in a constant reproductive number at a much lower level of transmission.
Figure 5: Estimated time-varying reproductive numbers (top) and posterior predictive epidemic trajectory (bottom) for both cities. Solid lines represent the means and the shaded regions represent the 95% credible intervals. Estimates for Wave 1 are shown in red and estimates for Wave 2 are shown in blue.
5 Discussion
In this work we have extended Bayesian behavioral change epidemic models in three important ways to make them more realistic and accurate. We expanded the SIR compartmental structure to incorporate additional data sources on deaths and hospitalizations in an SIHRD models. To the best of our knowledge, this is the first instance of an expanded compartmental structure in a fully stochastic population-averaged model, where both the parameters associated with transition probabilities and all transitions themselves are given distributions and are imputed when not observed. We investigated the incorporation of undetected infections by assuming cases are observed according to a constant detection probability and the true number of infections is estimated using data-augmented MCMC. Finally, we extended the behavioral change framework of Ward et al., (2023) to allow population alarm to be influence by multiple data sources. The proposed model specification also allows us to quantify the relative importance of each data source, so it can be used to investigate potential changes in importance over time.
Our simulation study illustrate several important features of fully Bayesian behavioral change models. First, the incorporation of undetected infections had a minimal impact on the estimation of the effective reproductive number. We believe this is primarily due to the large population size in our simulations, and that this incorporation would be more impactful when modeling epidemics in smaller populations. We did not evaluate this explicitly as our motivation was developing more accurate COVID-19 models in large populations, such as cities like Miami and Montréal. Secondly, we showed that the SIHRD model offers better estimation of $\mathcal{R}_{0}$ and the relative importance parameter in the multivariable alarm, illustrating the increased model complexity is beneficial. Finally when deaths are highly important in influencing behavior through the population alarm, it is critical that deaths are incorporated into the alarm function, particularly for estimating the reproductive number towards the end of the epidemic. However, when deaths are equally as important as cases or minimally important, both the estimation of $\mathcal{R}_{0}$ and overall model fit are comparable between an alarm based on cases only or the more flexible multivariable alarm.
Through analysis of multiple waves of COVID-19 in Miami and Montréal, we illustrate the benefit of our modeling approach in practice. Ignoring behavioral change completely resulted in much poorer model fit than the behavioral change models. We also found that cases were highly important in informing population behavior during the first wave of COVID-19 in both cities, but during the second wave in Miami, deaths were the primary influencing factor of behavior. This finding offers important insight into human behavioral change in response to a multiwave epidemic and may be due to perception going into Wave 2 that the virus was not as deadly as previously thought (Williams et al.,, 2021). The WAIC results from these analyses also indicated that there was minimal downside to using the more complex multivariable alarm function when only cases were important, and a large upside to including deaths when deaths were important. This is meaningful for using these models in practice as we don’t know which data source is influencing behavior prior to model fitting.
In this work we have considered a one-parameter growth function to describe the population alarm, but other functional forms could also be used. Functions used to describe the alarm must satisfy the following properties: for all $x$, $f(x)\in[0,1]$, $f(0)=0$, and for $x\leq y$, $f(x)\leq f(y)$. In words, the alarm can only take values between 0 and 1, must be 0 when there is no disease in the population, and is monotonically increasing as disease incidence increases. In addition to the one-parameter growth function we considered here, Ward et al., (2023) considered two other parametric functions which satisfy this criteria. A two-parameter constant change point model $f(x)=\delta\mathds{1}(x>H)$ with $H\in(0,\text{max}(x))$ $\delta\in[0,1]$, and a three-parameter modified Hill equation (Gesztelyi et al.,, 2012) $f(x)=\frac{\delta}{1+(x_{0}/x)^{\nu}}$ with $\delta\in[0,1]$, $x_{0}\in(0,\text{max}(x))$, and $\nu>0$ which can have a logistic looking shape to represent a smoothed version of the change point function. In considering multivariable extensions of these parametric functions, one must carefully think about identifiability of the overall alarm at any given time point. Both the change point and Hill functions rely on parameters estimating an asymptote ($\delta$) and a mid-point ($H$ or $x_{0}$) indicating the value of the observed data where the alarm is increasing. While this allows these functions to be more flexible, these parameterizations lead to identifiability issues when considering multiple data sources as the same epidemic could occur under various specifications of when the alarm components associated with cases or deaths “went off” (reaches the asymptote) and what level the alarm reaches based on each source (value of the asymptote). For example, an alarm fully attributed to cases may go off once smoothed incidence reaches 100 cases per day, but could also be fully attributed to deaths which reached three deaths at the same time. Or the alarm could be equally attributed to cases and death reaching these thresholds. This is problematic as all three scenarios could lead to the same alarm value as a function of epidemic time.
Alternatively, one may consider non-parametric approaches, such as a multivariate Gaussian process or combination of spline functions. However, care must be taken to ensure the final alarm is constrained between 0 and 1, which can be complicated when using multiple inputs. Additionally, one advantage of the current approach is the interpretability of the $\alpha$ parameter as the relative importance of one data source informing population alarm and its not obvious how to translate this interpretation to a non-parametric approach. Another possible approach would be the use of Bayesian model averaging (Hoeting et al.,, 1999), where each model contains a more complex alarm function based on a single data source and the posterior probabilities of each model would provide the relative importance of the data sources. We plan to pursue both the non-parametric and model averaging approaches as future work to extend this methodology.
Our SIR and SIHRD models have a few limitations that warrant discussion. First, they assume individuals are immediately infectious. While an Exposed compartment can be added to the behavioral change framework as in Ward et al., (2023), this requires strong priors on the latent period and additional data augmentation. Because the SIHRD model already augments several transitions and undetected infections, we omitted the latent period to reduce computational complexity. Second, we do not allow transmission from hospitalized individuals. Although a hospital transmission term could be added to the transmission probability, such effects are typically small, likely unidentifiable from available data, and would complicate interpretation of the alarm function. Third, we assume exponentially distributed infectious and hospitalization periods for the SIHRD model. More flexible distributions could be incorporated using the ‘linear chain trick’ to convolve multiple exponential distributions (Cushing,, 1977), however, this increase in model complexity would likely not have a substantial change on the inference on the reproductive number given that strong priors would be required.
Software
Software in the form of R code, together with data used and complete documentation is available at https://github.com/ceward18/multipleDataBCM.
Acknowledgments
Funding for the project was provided by the Canadian Statistical Sciences Institute (CANSSI) Distinguished Postdoctoral Fellowship. A. M. Schmidt and R. Deardon acknowledge financial support from the Natural Sciences and Engineering Research Council (NSERC) of Canada Discovery Grants program (Schmidt—RGPIN/2024-04312, Deardon—RGPIN/03292-2022). The authors also acknowledge the Minnesota Supercomputing Institute (MSI) at the University of Minnesota for providing resources that contributed to the research results reported within this paper. URL: http://www.msi.umn.edu.